# BondBERT: What we learn when assigning sentiment in the bond market

### v1

@@ 5 Results
For evaluation, we adopted an event-based methodology to isolate impactful news. Sentiment shocks were defined as daily sentiment values in the top or bottom 10th percentile. When multiple shocks occurred on the same day, their scores were averaged. We chose to evaluate the two sentiment tools using article topics related to the UK bond market, such as “UK inflation”, “UK bond market”, and “UK economic outlook”, as shown in Figure 3. The figures present the Pearson correlations between sentiment scores and UK bond returns across six economic topics, using a 10% daily sentiment shock, meaning using 10% of most negative and 10% of the most positive scores and a 7-day rolling window.
As we see, neither FinBERT nor BondBERT in Figures 3(a) and 3(b) achieved consistent high correlations amongst all topics and bond instruments. However, we do observe clear directional differences. FinBERT correlations are predominantly negative or near zero (i.e., between -0.11 and 0.05), consistent with the more general financial data on which it is trained, where positive macroeconomic news can imply falling bond prices. In contrast, BondBERT correlations are consistently positive across topics, with higher values, although it does show negative correlations for some instrument and topic combinations, similar to FinBERT. Moreover, FinGPT correlations in Figure 3(c) are mixed, with values clustering around zero and only limited alignment with bond price movements. This indicates that while FinGPT can capture some directional sentiment, its general-purpose training in finance limits performance on bond-specific topics. For Instruct-FinGPT in Figure 3(d), we observe that its correlations with bond returns are more stable and positive than those of FinGPT, and closer to the behaviour of BondBERT, so instruction tuning does seem to improve alignment with bond price movements. BondBERT, however, still achieves the strongest and most consistent correlations across topics.
Although the correlation magnitudes remain moderate, this sign reversal indicates BondBERT’s improved alignment of sentiment with bond price movements. In practical terms, BondBERT’s sentiment scores follow bond price direction more closely (i.e., positive sentiment with rising prices, negative sentiment with falling prices). This provides preliminary evidence of the suitability of BondBERT sentiment analysis of bond markets. Notably, FinBERT performs slightly better for the broad “UK Economic Outlook” topic, but BondBERT performs better across most other article categories.
Furthermore, we report the directional accuracy of sentiment signals for BondBERT, FinBERT, FinGPT, and Instruct-FinGPT relative to a 50% random baseline in Figure 4. Directional accuracy measures whether sentiment correctly predicts the sign of next-day bond returns following the same 10% daily sentiment shock.
Overall, accuracy levels for FinBERT and BondBERT are moderate, ranging from approximately 49% to 58% across the five different topics. BondBERT achieves its highest accuracy on the UK Bond Market and UK Inflation topics (58.14% and 57.58%, respectively), while FinBERT performs best on the UK Bond Market topic (56.52%). FinGPT and Instruct-FinGPT perform the worst, where all topics are below 40%. Notably, BondBERT outperforms the rest of the models on all topics. Given the low volatility of sovereign bonds, achieving even modest improvements in directional accuracy is challenging, since daily price changes are typically small, and random fluctuations can obscure sentiment trends.
These results suggest that while none of the models achieves strong predictive power when compared as such, BondBERT shows a more consistent advantage in aligning sentiment direction with subsequent bond price movements.
Table 3: Average LSTM results for the models across all bonds.
Metric BondBERT FinBERT FinGPT Instruct-FinGPT
Normalised RMSE 0.0079 0.00855 0.009 0.01
Information Coefficient 0.8 0.745 0.747 0.67
@@
Table 4: DM test p-values for BondBERT against the baselines across all bonds. Values in bold denote significance at the 5% level
Bond FinBERT FinGPT Instruct-FinGPT
@@
Additionally, we construct a pipeline to evaluate the predictive power of BondBERT versus FinBERT, FinGPT, and Instruct-FinGPT. We employ a LSTM model to predict next-day bond prices, using fixed-length rolling histories of daily bond transaction prices and the corresponding daily sentiment scores from each tool. For training, the model uses sequences based on an optimised rolling window and predicts the next-day price. All input features and targets are standardised to zero mean and unit variance.
We optimise the network using the Adam optimiser with weight decay and mean squared error (MSE) loss, applying early stopping based on validation loss to avoid overfitting. Hyperparameters such as hidden layer size, number of recurrent layers, dropout, learning rate, weight decay, and history length) are optimised using Optuna’s Bayesian sampler.
Model evaluation is conducted on the unseen test set using normalised RMSE and the information coefficient, with additional Diebold–Mariano (DM) tests (Newey–West corrected) to statistically assess differences between the four sentiment models. The results in Table 3 show a good predictive performance across models, with all achieving low RMSE and moderate to high information coefficient values. BondBERT achieves the lowest normalised RMSE (0.0079) and the highest information coefficient (0.8), suggesting it ranks price movements more accurately than the baselines.
We test differences in predictive accuracy using the DM test on normalised RMSE values, since forecast errors are defined at the observation level and allow for statistical comparison. However, information coefficient is computed as a cross-sectional correlation across only ten bonds, providing too few observations for meaningful inference. Thus, the information coefficient values are reported descriptively rather than tested for significance. The null hypothesis of the DM test is that competing forecasts have the same population accuracy. BondBERT rejects the null hypothesis in half of the bonds, being showing statistically significant improvements in normalised RMSE at the $0.05$ level compared to the baseline models, as seen in Table 4.
In summary, we notice that BondBERT reverses the negative bias of FinBERT, FinGPT, and Instruct-FinGPT, achieving consistently positive correlations. It also delivers higher directional accuracy despite the low volatility of sovereign bonds, and improves LSTM forecasts, statistically significantly different over the baselines.
@@ 6 Conclusions and Future work
This study introduces BondBERT, a bond-specific sentiment analysis model, and demonstrates its advantage over three other sentiment tools in aligning sentiment with bond price movements. We showed the importance of a transferable strategy for adapting LLMs to markets where sentiment dynamics diverge from equity assumptions, easily reproducible, and generalisable to other non-equity classes, such as commodities, derivatives, and real estate. Our results show that BondBERT reverses the wrong-sign bias observed in equity-trained models and delivers more consistent forecasting improvements. The analysis is restricted to UK sovereign bonds, since the illiquidity of corporate bonds prevents a meaningful evaluation in that segment, and the current model is trained on synthetic sentiment labels generated by GPT-4.1-nano. While this approach is cost-effective and scalable, it may not fully capture domain expertise, and thus constrains the fidelity of the training data. Thus, we plan to introduce human annotation and semi-supervised refinement to improve the quality of sentiment labels beyond synthetic generation. Moreover, we aim to expand the dataset in both temporal and market scope, incorporating international sovereign and corporate bonds, with higher liquidity. These extensions will enable testing the generalisability of BondBERT and strengthen its potential applications in trading, risk management, and policy. While proprietary bond price data restricts direct replication, we aim to release our labelling pipeline, prompts, and model to enable reproduction of core experiments. Beyond finance, we also highlight a broader AI challenge, including learning under weak effect sizes and noisy supervision, where sentiment often diverges from intuitive human interpretations. While BondBERT is not a standalone agent, it demonstrates how specialised perception modules can be embedded into intelligent decision-support agents for fixed income markets. Such modules could interact with forecasting and trading agents as part of multi-agent systems, extending the role of NLP in financial AI.
In the longer term, the goal is to incorporate predictive models into an agentic-AI framework to provide bespoke investment advice to individual investors via a mobile application. For more details, see the artificial intelligence for collective intelligence (AI4CI) vision, described in Bullock et al., (2024).
@@ ACKNOWLEDGEMENTS
This work was supported by UK Research and Innovation (UKRI) Engineering and Physical Sciences Research Council (EPSRC) Grant Number EP/Y028392/1: AI for Collective Intelligence (AI4CI). TB was supported by a Faculty of Engineering Summer Research Internship. The authors thank Vincent Grandjean and Oliver Haste at Propellant Digital for access to data and for their expert advice and support.

### v2

@@ 5 Results
For evaluation, we adopted an event-based methodology to isolate impactful news. Sentiment shocks were defined as daily sentiment values in the top or bottom 10th percentile. When multiple shocks occurred on the same day, their scores were averaged.
Figure 2 presents Pearson correlations between sentiment scores and UK bond returns on six economic topics, using a 10% daily sentiment shock (i.e., 10% most negative and 10% most positive scores) and a 7-day rolling window. Overall, we see that BondBERT has positive correlations across most topics—with negative correlations only for topic “UK Bond Market”. FinBERT only has positive correlations for the topic “UK Economic Outlook”—the other topics are neutral or negative—consistent with the general financial data on which it is trained. FinGPT exhibits positive and negative correlations, which are relatively strong and more consistent than FinBERT. Finally, Instruct-FinGPT displays strong negative correlations for topic “UK Bond Market” and weaker, less consistent correlations elsewhere. In general, BondBERT shows the best alignment between bond sentiment and bond price movements, which provides preliminary evidence of the suitability of BondBERT for sentiment analysis of bond markets.
Figure 3 compares the directional accuracy of the four models, relative to a 50% random baseline (horizontal red line). Directional accuracy is measured as the directional prediction (up/down) of the next-day bond returns after a 10% daily sentiment shock. We see that BondBERT (blue bars) has the highest directional accuracy on all five topics and is the only model that consistently outperforms the random baseline. FinBERT (orange bars) performs second best, but is inconsistent and does not outperform the random baseline on two of the five topics. FinGPT (green bars) and Instruct-FinGPT (red bars) both perform poorly, well below the random baseline on every topic. These results clearly show that BondBERT is the best performing model, and the only model with directional accuracy better than random.
Figure 3: Model performance comparison, showing directional accuracy of sentiment relative to bond returns.
To assess the predictive power of the sentiment models, we used an LSTM to predict next-day bond prices using the fixed-length rolling histories of daily bond transaction prices and the daily sentiment scores from each model as input. All input features and targets were standardised to zero mean and unit variance. The LSTM was optimised using the Adam optimiser, with weight decay and mean squared error (MSE) loss, applying early stopping based on validation loss to avoid overfitting. Hyperparameters such as hidden layer size, number of recurrent layers, dropout, learning rate, weight decay, and history length were optimised using Optuna’s Bayesian sampler. Model evaluation is conducted on the unseen test set using normalised RMSE and the information coefficient (IC), and Diebold–Mariano (DM) tests with Newey-West corrected errors to assess whether forecast accuracy differences are statistically significant.
Table 2 presents the LSTM results, showing good predictive performance in all models, with low RMSE and moderate to high IC values. BondBERT achieves the lowest normalised RMSE and the highest IC, which suggest that it can be used to predict price movements more accurately than the other models. Table 3 presents the p-values of the DM test on normalised RMSE values of BondBERT against the other three models. The null hypothesis of the DM test is that competing forecasts between BondBERT and the competing model have the same population accuracy. We see that the null is rejected in favour of BondBERT in half of the bonds (values in bold), which demonstrates that BondBERT generates statistically significant improvements in normalised RMSE.
In summary, we have shown that BondBERT is able to achieve consistently positive sentiment correlations, which reverses the negative bias exhibited by FinBERT, FinGPT, and Instruct-FinGPT. We have also shown that BondBERT has higher directional accuracy and significantly outperforms other models on next-day price prediction.
Table 2: Mean LSTM results per model, showing Normalised RMSE and Information Coefficient (IC).
Metric BondBERT FinBERT FinGPT Inst-FinGPT
RMSE 0.0079 0.0086 0.009 0.01
IC 0.8000 0.7450 0.747 0.67
@@
Table 3: DM test p-values for BondBERT against the baselines. Values in bold denote significance at the 5% level.
Instrument ID FinBERT FinGPT Inst-FinGPT
@@ 6 Conclusions and Future work
This study introduces BondBERT, a bond-specific sentiment model, and demonstrates its advantage over three other tools in aligning sentiment with bond price movements. We showed the importance of a transferable strategy for adapting LLMs to markets where sentiment dynamics diverge from equity assumptions and generalise to other asset classes. BondBERT reverses the wrong-sign bias seen in equity-trained models and delivers more consistent forecasting improvements. The analysis is restricted to UK sovereign bonds, since the illiquidity of corporate bonds prevents a meaningful evaluation, and the model is trained on synthetic labels generated by GPT-4.1-nano. Although this approach is cost-effective and scalable, it may not fully capture domain expertise and constrains label fidelity. We plan to introduce human annotation and semi-supervised refinement to improve label quality. Moreover, our goal is to expand the dataset across, incorporating international sovereign and corporate bonds with higher liquidity. These extensions will test the generalisability of BondBERT and strengthen its potential applications in trading, risk management, and policy.
Beyond finance, we highlight a broader AI challenge, learning under weak effect sizes and noisy supervision, where sentiment diverges from intuitive human interpretations. While BondBERT is not a standalone agent, it demonstrates how specialised perception modules can be embedded into decision-support agents for fixed income markets and interact with forecasting and trading agents as part of multi-agent systems, extending the role of NLP in financial AI. In the longer term, the goal is to incorporate predictive models into an agentic-AI framework to provide bespoke investment advice to individual investors via a mobile application. For more details, see the artificial intelligence for collective intelligence (AI4CI) vision, described in Bullock et al., (2024).
@@ ACKNOWLEDGEMENTS
This work was supported by UK Research and Innovation (UKRI) Engineering and Physical Sciences Research Council (EPSRC) [grant number EP/Y028392/1]: AI for Collective Intelligence (AI4CI). Toby Barter was supported by a Faculty of Engineering Summer Research Internship. The authors thank Vincent Grandjean and Oliver Haste at Propellant Digital for expert advice and data access.